OpenStudy (anonymous):

Given the linear equation y = 5x – 1, find the y-coordinates of the points (-1, ), (2, ), and (1, ). Please show all of your work.

OpenStudy (anonymous):

Just put x into the equation and you get y values. For the first one, it'd be (-1, -6)

OpenStudy (anonymous):

was the rest of them i dnt know this stuff

OpenStudy (anonymous):

did yu multiply or add the x with the y

OpenStudy (anonymous):

like 5*2=10

OpenStudy (anonymous):

Just put your x values, -1, 2 and 1 into the equation y = 5x -1, and you get a y value.

OpenStudy (anonymous):

Your co-ordinates are in the form (x, y)

OpenStudy (anonymous):

i dnt get it

OpenStudy (anonymous):

Ok, you see your co-ordinates that you have, (-1, y) (2, y) and (1, y). These give you 3 x values. Put these into the equation to get a y value. For example, with x = -1, You put it in to the equation and you get y=5(-1) -1 = -5 - 1 = -6, so y=-6, and your co-ordinate is (-1, -6)

OpenStudy (anonymous):

gud thang this has nthn 2 do with money cuss i'd be gettin beat

OpenStudy (anonymous):

and wut will be the x value if the y value is -1,-6

OpenStudy (anonymous):

Well to find x, you need to rearrange the formula for x. y= 5x - 1 y + 1 = 5x (y + 1)/5 = x putting -1 in, we would get 0/5 = x, so x = 0 Putting -6 in, we get -5/5 = x, so x = -1

OpenStudy (anonymous):

THANKS YU WERE AND ARE ALOT OF HELP YU DESERVE THE MEDALS IM GIVIN YU